# To Supervise or Not to Supervise: Understanding and Addressing the Key Challenges of 3D Transfer Learning

### v1

@@ Abstract
Transfer learning has long been a key factor in the advancement of many fields including 2D image analysis. Unfortunately, its applicability in 3D data processing has been relatively limited. While several approaches for 3D transfer learning have been proposed in recent literature, with contrastive learning gaining particular prominence, most existing methods in this domain have only been studied and evaluated in limited scenarios. Most importantly, there is currently a lack of principled understanding of both when and why 3D transfer learning methods are applicable. Remarkably, even the applicability of standard supervised pre-training is poorly understood. In this work, we conduct the first in-depth quantitative and qualitative investigation of supervised and contrastive pre-training strategies and their utility in downstream 3D tasks. We demonstrate that layer-wise analysis of learned features provides significant insight into the downstream utility of trained networks. Informed by this analysis, we propose a simple geometric regularization strategy, which improves the transferability of supervised pre-training. Our work thus sheds light onto both the specific challenges of 3D transfer learning, as well as strategies to overcome them.
@@ 1 Introduction
Figure 1: Analyzing and improving 3D transfer learning. In this work, we perform the first in-depth study of the different components of network pre-training that influence the outcome of 3D transfer learning (components in solid orange). This includes the source data domain in relation to the downstream target data, the choice of architecture, the importance of early vs later layers, and the pre-training design choices. We also show that improvements to the pipeline can be achieved through regularization of the early layers, by promoting the prediction of geometric properties.
@@
In spite of these challenges, several works have recently emerged aiming to enable transfer learning in 3D, especially by focusing on new pre-training strategies for point clouds, e.g., [50, 21, 48]. Among these strategies, several approaches have shown that self-supervised strategies (e.g. contrastive learning, point cloud reconstruction) can improve performance on different downstream data (e.g. shapes, scenes) and on select set of tasks and datasets.
Despite the growing number of these approaches, unfortunately, most existing methods are only presented and evaluated in very specific settings, and missing comparison to baselines such as supervised pre-training. Specifically, there are several questions not addressed in existing literature, including: 1) The utility of supervised compared to contrastive pre-training, 2) The role of the architecture’s inherent properties in the success of 3D transfer learning, and 3) The impact of regularization strategies on the efficacy of 3D pre-training.
By performing the first in-depth investigation of the most prominent approaches, we observe that in the linear probing setting, supervised pre-training leads to superior performance compared to contrastive learning across almost all tested architectures, even under relatively significant domain shift. This suggests that within 3D deep learning, contrastive learning does not always lead to more “universally useful” features. At the same time, we also observe that in the full fine-tuning scenario, contrastive pre-training often strongly outperforms supervised learning. By analyzing gradients of pre-trained networks, we show that the early layers learned with supervised pre-training are not easily adaptable across new datasets, compared to those learned with contrastive learning. This sheds new and specific light on the differences between these pre-training strategies in the context of 3D shape analysis. Finally, we demonstrate that a simple regularization scheme applied to the early layers during supervised pre-training can overcome this limitation and leads to networks that can be fine-tuned efficiently.
@@
    We perform the first comparison, within a single consistent evaluation framework, of supervised and contrastive pre-trained models using different 3D backbone architectures, assessing their transfer learning performance on most common downstream data and tasks. We separate between the linear probing and fine-tuning settings and analyze the difference in performance.
@@ 2 Related Work
Finally, we note that several works have aimed to analyze the factors behind successful transfer learning approaches in 2D image tasks [34, 56, 30]. These works point to the special importance of early and mid-level features and to layer-wise analysis of learned representations. Such analysis has not yet been performed for 3D data, and our main objective is to fill this gap by performing a first comprehensive investigation of this topic. As we mention below, our work highlights the special nature of learning on 3D data, thus shedding light on the challenges of 3D transfer learning and possible ways to address them.
@@ 3.1 Datasets
Pre-training Dataset.  Prior 3D transfer learning approaches [45, 1, 50, 21] have typically focused on pre-training datasets with limited domain shifts to downstream tasks, such as within the synthetic shape setting [5, 47] or the real scene scenario [9, 3]. Interestingly, among them, PointContrast [50] argued against pre-training on synthetic data and stated that supervised pre-training is an upper bound to performance (see Sec. 3.1 therein). Our experimental findings (Sec. 5), however, suggest that this is not always the case and help to refine this statement. We thus select ShapeNet [5] to evaluate the utility of pre-training with synthetic data, and focus on its transfer utility in downstream tasks with varying domain shifts, as elaborated below.
@@
  • ModelNet40 [47] is a synthetic object-level dataset with 3D shapes resembling those in ShapeNet. It consists of 12,311 shapes and 40 object classes for the shape classification task, 20 of these classes can be found in ShapeNet.
  • ScanObjectNN [44] is a real object-level dataset with scanned 3D shapes featuring noise and background, vastly different from the clean shapes in ShapeNet. There are 15,000 shapes and 15 object classes for the shape classification task, 9 of these classes can be found in ShapeNet.
@@ 4 Pre-training Approaches
As mentioned above, in our analysis, we concentrate on comparing supervised and contrastive learning approaches. Although our primary focus is on these pre-training strategies, we include additional results to alternative methods (including shape-level contrastive learning and a reconstruction-based approach PointDAE [54]) in the supplementary materials.
@@
where $y_{i}$ is the ground truth label, and $p_{i}$ is the predicted probability of each class $i$ for an input shape.
Contrastive pre-training is employed for comparisons with supervised pre-training. We use point-level contrastive learning, based on PointContrast [50]. Specifically, for contrastive pre-training, we first introduce a set of invariances by applying data augmentation. We start by normalizing a point cloud and perform random geometric transformations, including translation with magnitude of 0.5, scaling between 80% and 125% and rotation of magnitude 45. We also simulate partial data by cropping 50% the point cloud. More details about transformations can be found in the supplementary materials. By combining these transformations, we start from an input point cloud and generate two augmented views $P_{i}$ and $P_{j}.$
@@ 5.2 Results
Unique architectural properties.  While supervised pre-training generally leads to useful final layer features, the results significantly depend on the choice of architectures. Specifically, we observe that the hierarchical nature of DGCNN and PointMLP results in more informative and general features compared to global architectures such as PointNet. In additon, the transformer architecture (PCT) prefers the contrastive approach, particularly under strong domain shift.
Table 1: Evaluation on S3DIS semantic segmentation task of pre-trained models. IoU metric used. Random init is a randomly initialized model.
@@
Transfer learning with fine-tuning.  We observe a notable shift of behavior when performing full fine-tuning, compared to linear probing, as shown in Figures 2(a)2(b) (dashed bars). We remark that for several architectures, such as DGCNN, MinkowskiNet and PCT, contrastive pre-training leads to better results compared to supervised learning under full fine-tuning. This suggests that the utility of pre-trained models highly depends on how those models are used in the downstream tasks. Specifically, we note that the utility of final layer features (such as with linear probing) is not always a good predictor for the performance under fine-tuning, as shown, e.g., in the case of DGCNN. In Section 4 we analyze this behavior in depth and observe the critical role of early layers and their ability to adapt to downstream data. Our analysis also highlights that the types of solutions obtained with supervised and contrastive pre-training are different, making a distinction between generalizable and adaptable features.
In addition, we note that for most architectures, including MinkowskiNet, supervised pre-training is not an “upper bound” on the fine-tuning performance, differently from the claim made in [50], and that other pre-training strategies can lead to better results.
Unique architectural properties.  PointNet, with its global feature aggregation strategy, tends to overfit to the source training dataset, and is thus not able to extract information that is useful for fine-tuning. MinkowskiNet’s performance is influenced by variations in point cloud density and scale. Its ability to process large-scale scenes explains its high semantic segmentation performance (Table 1), but highlights its dependency on the sampling density for effective transfer learning (i.e. there is improvement from small-scale synthetic shapes to natural noisy scanned shapes, but not to large-scale scenes). In contrast, PointMLP benefits more from supervised pre-training when the downstream task is classification (e.g., ModelNet40 and ScanObjectNN) but still attains its limits in the more challenging semantic segmentation task.
@@ 6 Feature Analysis
Discriminative capability of early layers.  To illustrate the behavior of the early layers, we plot a t-SNE visualization of the first-layer features of pre-trained models in Figure 3 on the ModelNet10 dataset. We observe the appearance of discernible clusters even in these early layers and without any fine-tuning. This clustering in the feature space suggests that early layers possess a class-discriminative capacity. This points to the potential of early layers to contribute significantly to the downstream task, unlike in the 2D domain where this property is rarely observed.
@@
Gradient norm of layer weights.  To understand how the models behave in the fine-tuning setting, we analyze the norm of the gradients of pre-trained networks with respect to downstream data. Figure 4 shows the layer-wise gradient norm, computed using a cross-entropy loss on the downstream data using frozen pre-trained weights. We observed that for contrastive pre-training, early layers exhibit a higher gradient norm compared to the last layer. This finding further substantiates the need, not only for last layers, but also for early layers to adapt to new data/tasks when dealing with 3D data for successful transfer learning.
@@ 7.1 Normal Prediction
Our analysis above has revealed various challenges in pre-training that can hinder effective transfer learning. Specifically, we have observed a correlation between low gradient norms in early layer weights and diminished fine-tuning performance, indicating a decreased adaptability of the learned features when moving from source to downstream tasks.
@@
We apply this regularization only to a set of early layers, whereas the output features of the entire architecture continue to serve the primary pre-text task. This strategy infuses purely geometric information into early layers with the goal of making them universally applicable and reduce the dependency to the pre-text task. Let $H_{\text{normal}}(\mathbf{f}_{l})$ denote the normal prediction head, $\mathbf{f}_{l}$ the point-wise features of layer $l$, and $\mathbf{n}$ the ground truth normals. The combined training objective comprises the pre-training loss $L_{\text{pre-train}}$ and the regularization loss $L_{\text{regul}}$. We use absolute cosine similarity in $L_{\text{regul}}$ since estimated normals are not oriented. Our total pre-training loss can be formulated as:
@@ 7.2 Evaluation
The experimental setup for pre-training and fine-tuning remains consistent with our previous evaluation. We use normals computed through PCA on local neighborhoods comprising 30 points within point clouds of size 2048. We explore the optimal number of layers required for the most effective regularization, and find that regularizing the features of the first two layers yields the best results. The subsequent results and analysis focus on the optimal setting for DGCNN, while other layer choices and architectures are studied in the supplementary.
@@ 8 Summary and Key Takeaway Remarks
Second, we observe that unlike the 2D case, features learned even in the earliest layers of 3D architectures tend to be class-specific, which can hinder their utility in transfer learning. This highlights the importance of adaptability of early layers for 3D transfer learning.
@@ 9 Conclusion, Limitations and Future Work
In this paper, we have conducted the first in-depth qualitative and quantitative investigation of the key factors when performing 3D transfer learning with supervised vs contrastive pre-training and proposed a regularization approach informed this analysis. We evaluated several architectures under pre-training approaches and found that their performance changes with the transfer learning protocol. We also examined early layer importance through their discriminative capability and fine-tuning adaptability, shedding light on their utility and importance for 3D transfer learning. Finally, we correlated this analysis with a new regularization approach that targets early layers to improve on downstream performance. Overall, our work establishes a consistent evaluation framework, presents detailed analysis tools, and proposes an appropriate simple method for successful transfer learning on 3D data, which can both inform and allow to compare future designs.
As our main focus was on the careful analysis of the factors that contribute to successful transfer learning in 3D, we did not investigate all possible combinations of the source datasets, architectures, pre-training strategies, fine-tuning approaches, and downstream tasks. The resulting combinatorial complexity would incur very significant computational costs, and furthermore might obscure the possible analytical insights. Nevertheless, in the future, an analysis of other tasks and pre-training strategies can fit within and complement the framework that we established. Finally, as suggested by our analysis, there is currently a need for new architectures and pre-training approaches, that are robust under changes of sampling density and can lead to multi-scale features, which would generalize and adapt to new downstream data. Exploring the possible solution space is an exciting direction for future work.

### v2

# To Supervise or Not to Supervise: Understanding and Addressing the Key Challenges of Point Cloud Transfer Learning

@@ Abstract
Transfer learning has long been a key factor in the advancement of many fields including 2D image analysis. Unfortunately, its applicability in 3D data processing has been relatively limited. While several approaches for point cloud transfer learning have been proposed in recent literature, with contrastive learning gaining particular prominence, most existing methods in this domain have only been studied and evaluated in limited scenarios. Most importantly, there is currently a lack of principled understanding of both when and why point cloud transfer learning methods are applicable. Remarkably, even the applicability of standard supervised pre-training is poorly understood. In this work, we conduct the first in-depth quantitative and qualitative investigation of supervised and contrastive pre-training strategies and their utility in downstream 3D tasks. We demonstrate that layer-wise analysis of learned features provides significant insight into the downstream utility of trained networks. Informed by this analysis, we propose a simple geometric regularization strategy, which improves the transferability of supervised pre-training. Our work thus sheds light onto both the specific challenges of point cloud transfer learning, as well as strategies to overcome them.
@@ 1 Introduction
Figure 1: Analyzing and improving point cloud transfer learning. In this work, we perform the first in-depth study of the different components of network pre-training that influence the outcome of point cloud transfer learning (components in solid orange). This includes the source data domain in relation to the downstream target data, the choice of architecture, the importance of early vs later layers, and the pre-training design choices. We also show that improvements to the pipeline can be achieved through regularization of the early layers, by promoting the prediction of geometric properties.
@@
In spite of these challenges, several works have recently emerged aiming to enable transfer learning for point clouds, especially by focusing on new pre-training strategies, e.g., [53, 21, 51]. Among these strategies, several approaches have shown that self-supervised strategies (e.g. contrastive learning, point cloud reconstruction) can improve performance on different downstream data (e.g. shapes, scenes) and on select set of tasks and datasets.
Despite the growing number of these approaches, unfortunately, most existing methods are only presented and evaluated in very specific settings, and missing comparison to baselines such as supervised pre-training. Specifically, there are several questions not addressed in existing literature, including: 1) The utility of supervised compared to contrastive pre-training, 2) The role of the architecture’s inherent properties in the success of point cloud transfer learning, and 3) The impact of regularization strategies on the efficacy of point cloud pre-training.
By performing the first in-depth investigation of the most prominent approaches, we observe that in the linear probing setting, supervised pre-training leads to superior performance compared to contrastive learning across almost all tested architectures, even under relatively significant domain shift. This suggests that within 3D deep learning, contrastive learning does not always lead to more “universally useful” features. At the same time, we also observe that in the full fine-tuning scenario, contrastive pre-training can outperform supervised learning. By analyzing gradients of pre-trained networks, we show that the early layers learned with supervised pre-training are not easily adaptable across new datasets, compared to those learned with contrastive learning. This sheds new and specific light on the differences between these pre-training strategies in the context of point cloud analysis. Finally, we demonstrate that a simple regularization scheme applied to the early layers during supervised pre-training can overcome this limitation and leads to networks that can be fine-tuned efficiently.
@@
    We perform the first comparison, within a single consistent evaluation framework, of supervised and contrastive pre-trained models using different point cloud backbones, assessing their transfer learning performance on common downstream data and tasks. We separate between the linear probing and fine-tuning settings and analyze the difference in performance.
@@ 2 Related Work
Finally, we note that several works have aimed to analyze the factors behind successful transfer learning approaches in 2D image tasks [37, 59, 31]. These works point to the special importance of early and mid-level features and to layer-wise analysis of learned representations. Such analysis has not yet been performed for 3D data, and our main objective is to fill this gap by performing a first comprehensive investigation of this topic. As we mention below, our work highlights the special nature of learning on 3D data, thus shedding light on the challenges of point cloud transfer learning and possible ways to address them.
@@ 3.1 Datasets
Pre-training Dataset.  Prior 3D transfer learning approaches [48, 1, 53, 21] have typically focused on pre-training datasets with limited domain shifts to downstream tasks, such as within the synthetic shape setting [5, 50] or the real scene scenario [9, 3]. Interestingly, among them, PointContrast [53] argued against pre-training on synthetic data and stated that supervised pre-training is an upper bound to information gain from pre-training (see Sec. 3.1 therein). Our experimental findings (Sec. 5), however, suggest that this is not always the case and help to refine this statement. We thus select ShapeNet [5] to evaluate the utility of pre-training with synthetic data, and focus on its transfer utility in downstream tasks with varying domain shifts, as elaborated below.
@@
  • ModelNet40 [50] is a synthetic object-level dataset with 3D shapes resembling those in ShapeNet. It consists of 12,311 shapes and 40 object classes for the shape classification task, and 20 of these classes can be found in ShapeNet.
  • ScanObjectNN [47] is a real object-level dataset with scanned 3D shapes featuring noise and background, vastly different from the clean shapes in ShapeNet. There are 15,000 shapes and 15 object classes for the shape classification task, and 9 of these classes can be found in ShapeNet.
@@ 4 Pre-training Approaches
As mentioned above, in our analysis, we focus on comparing supervised and contrastive learning approaches. Although our primary focus is on these pre-training strategies, we include additional results to alternative methods (including shape-level contrastive learning and a reconstruction-based approach PointDAE [57]) in the supplementary materials.
@@
where $y_{i}$ is the ground truth label, and $p_{i}$ is the predicted probability of each class $i$ for an input shape. We also use data augmentation (translation, scale and rotation) to avoid the orientation bias introduced by the oriented shapes of ShapeNet, which are also present in the ModelNet40 downstream data.
Contrastive pre-training is employed for comparisons with supervised pre-training. We use point-level contrastive learning, based on PointContrast [53]. Specifically, for contrastive pre-training, we first introduce a set of invariances by applying data augmentation. We start by normalizing a point cloud and perform random geometric transformations, including translation with magnitude of 0.5, scaling between 80% and 125%, and rotation of magnitude 45. We also simulate partial data by cropping 50% the point cloud.
More details about transformations can be found in the supplementary materials. By combining these transformations, we start from an input point cloud and generate two augmented views $P_{i}$ and $P_{j}.$
@@ 5.2 Results
Unique architectural properties.  While supervised pre-training generally leads to useful final layer features, the results significantly depend on the choice of architectures. Specifically, we observe that the hierarchical nature of DGCNN and PointMLP results in more general features compared to global architectures such as PointNet. In addition, PCT’s transformer architecture is adept at capturing global dependencies which overfits less with unsupervised (contrastive) pre-training.
Table 1: Evaluation on S3DIS semantic segmentation fine-tuning of pre-trained models. IoU metric used. Random init is a randomly initialized model.
@@
Transfer learning with fine-tuning.  We observe a notable shift of behavior when performing full fine-tuning, compared to linear probing, as shown in Figures 2(a)2(b) (dashed bars). We remark that for several architectures, such as DGCNN, MinkowskiNet and PCT, contrastive pre-training leads to better results compared to supervised learning under full fine-tuning (with a $\sim$0.1accuracy advantage in the on-par scenarios). This suggests that the utility of pre-trained models highly depends on how those models are used in the downstream tasks. Specifically, we note that the utility of final layer features (such as with linear probing) is not always a good predictor for the performance under fine-tuning, as shown, e.g., in the case of DGCNN. In Section 4 we analyze this behavior in depth and observe the critical role of early layers and their ability to adapt to downstream data. Our analysis also highlights that the types of solutions obtained with supervised and contrastive pre-training are different, making a distinction between generalizable (features ready to be used in multi-task settings) and adaptable features.
In addition, we note that for most architectures, including MinkowskiNet, supervised pre-training is not an “upper bound” to information gain from pre-training, differently from the claim made in [53], and that other pre-training strategies can lead to better results.
Unique architectural properties.  PointNet, with its global feature aggregation strategy, tends to overfit to the source training dataset, and is thus not able to extract information that is useful for fine-tuning. MinkowskiNet’s performance is influenced by variations in point cloud density and scale. Its ability to process large-scale scenes explains its high semantic segmentation performance (Table 1), but highlights its dependence on the sampling density for effective transfer learning.
@@ 6 Feature Analysis
Discriminative capability of early layers.  To illustrate the behavior of the early layers, we plot a t-SNE visualization of the first-layer features of pre-trained models in Figure 3 on the ModelNet10 dataset. We observe the appearance of discernible clusters even in these early layers and without any fine-tuning. This clustering in the feature space suggests that early layers possess a class-discriminative capacity. This points to the potential of early layers to contribute significantly to the downstream task, unlike in the 2D domain where early layers are typically "universal" and do not need to be adapted [32].
@@
Gradient norm of layer weights.  To understand how the models behave in the fine-tuning setting, we analyze the norm of the gradients of pre-trained networks with respect to downstream data. Figure 4 shows the layer-wise gradient norm, computed using a cross-entropy loss on the downstream training data using frozen pre-trained weights. We observed that for contrastive pre-training, early layers exhibit a higher gradient norm compared to the last layer. This finding further substantiates the need, not only for last layers, but also for early layers to adapt to new data/tasks when dealing with 3D data for successful transfer learning.
@@ 7.1 Normal Prediction
The qualitative insights of our analysis have revealed various challenges in pre-training that can hinder effective transfer learning. Specifically, we have observed a correlation between low gradient norms in early layer weights and diminished fine-tuning performance, indicating a decreased adaptability of the learned features when moving from source to downstream tasks.
@@
We apply this regularization only to a set of early layers, whereas the output features of the entire architecture continue to serve the primary pre-text task. This strategy infuses purely geometric information into early layers with the goal of making them universally applicable and reducing the dependency to the pre-text task. Let $H_{\text{normal}}(\mathbf{f}_{l})$ denote the normal prediction head, $\mathbf{f}_{l}$ the point-wise features of layer $l$, and $\mathbf{n}$ the ground truth normals. The combined training objective comprises the pre-training loss $L_{\text{pre-train}}$ and the regularization loss $L_{\text{regul}}$. We use absolute cosine similarity in $L_{\text{regul}}$ since estimated normals are not oriented. Our total pre-training loss can be formulated as:
@@ 7.2 Evaluation
The experimental setup for pre-training and fine-tuning remains consistent with our previous evaluation. We use normals computed through PCA on local neighborhoods comprising 30 points within point clouds of size 2048. We explore the optimal number of layers required for the most effective regularization, and find that regularizing the features of the first two layers yields the best results. Through empirical testing to determine the optimal layers for regularization, we discovered that targeting early layers for regularization yields better outcomes than applying it to the encoder’s output layer alone. Ablation of layer choice and other architectures is included in the supplementary.
@@ 8 Summary and Key Takeaway Remarks
Second, we observe that unlike the 2D case, features learned even in the earliest layers of 3D architectures tend to be class-specific, which can hinder their utility in transfer learning. This highlights the importance of adaptability of early layers for point cloud transfer learning.
@@ 9 Conclusion, Limitations and Future Work
In this paper, we have conducted the first in-depth qualitative and quantitative investigation of the key factors when performing point cloud transfer learning with supervised vs contrastive pre-training and proposed a regularization approach informed by this analysis. We evaluated several architectures under pre-training approaches and found that their performance changes with the transfer learning protocol. We also examined early layer importance through their discriminative capability and fine-tuning adaptability, shedding light on their utility and importance for point cloud transfer learning. Finally, we correlated this analysis with a new regularization approach that targets early layers to improve on downstream performance. Overall, our work establishes a consistent evaluation framework, presents detailed analysis tools, and proposes an appropriate simple method for successful transfer learning on point cloud data, which can both inform and allow to compare future designs.
As our main focus was on the careful analysis of the factors that contribute to successful transfer learning for point clouds, we did not investigate all possible combinations of the source datasets, architectures, pre-training strategies, fine-tuning approaches, and downstream tasks. The resulting combinatorial complexity would incur very significant computational costs, and furthermore might obscure the possible analytical insights. Nevertheless, in the future, an analysis of other tasks and pre-training strategies can fit within and complement the framework that we established. Finally, as suggested by our analysis, there is currently a need for new architectures and pre-training approaches, that are robust under changes of sampling density and can lead to multi-scale features, which would generalize and adapt to new downstream data. Exploring the possible solution space is an exciting direction for future work.
Acknowledgements  Parts of this work were supported by the ERC Starting Grant 758800 (EXPROTEA), ERC Consolidator Grant 101087347 (VEGA), ANR AI Chair AIGRETTE, as well as gifts from Ansys and Adobe Research.